# Exploring Multi-Document Information Consolidation for Scientific Sentiment Summarization

### v1

@@ 4 Sentiment-Aware Evaluation on Information Consolidation
In this section, we focus more on how to evaluate sentiments of the generated summaries or meta-reviews in meta-review generation based on our proposed framework. We propose FacetEval and FusionEval which are reference-based and reference-free metrics, respectively.
@@ 4.1 Measuring Sentiment Similarity to Human-Written Meta-Review
To assess the quality of generated meta-reviews, we propose a reference-based evaluation metric, FacetEval, measuring the sentiment consistency $c$ between the generated meta-review and the corresponding human-written meta-review in all criteria facets. Different from the generic evaluation metrics for abstractive summarization or text generation which mostly adopt surface-form matching, we focus more on criteria facets and their corresponding sentiment levels in human-written and model-generated meta-reviews.
Specifically, we use the distribution of sentiments in all criteria facets to represent the meta-review and use the cosine similarity of the two vectors as the final score $s$.
@@
where $\bm{d}^{f}=[P^{+}_{f},P_{f},N^{+}_{f},N_{f},O_{f}]$. Results shown in Table 6 suggest that different reviews are consistent in the sentiment to Compliance while there is much lower consistency in Clarity and Novelty. Moreover, we find that conflict reviews [^11] The same as in PeerSum, if any two reviews have ratings where the gap is larger than 4 they are conflict reviews. would prefer showing conflicts in Advancement, Novelty, Clarity and Overall. This is also consistent with our typical understanding in peer reviews and occasional conflicts among them.
Criteria Facet w/ Conflicts w/o Conflicts Advancement 0.463 (0.135) 0.551 (0.137) Soundness 0.526 (0.158) 0.501 (0.110) Novelty 0.300 (0.159) 0.357 (0.168) Overall 0.433 (0.147) 0.597 (0.172) Clarity 0.317 (0.133) 0.337 (0.145) Compliance 0.827 (0.071) 0.771 (0.118)
@@
4.2 Measuring Sentiment Fusion for Individual Facets
Sentiments in the generated meta-reviews should be in line with the aggregate sentiment from the individual source documents. As GPT-4 has the potential to predict sentiments in the human-written meta-reviews from judgements in source reviews and discussions in Section 3.3, we present a reference-free evaluation metric, FusionEval, which assesses the consistency between the sentiments in the generated meta-review and that predicted by the zero-shot sentiment level prediction with GPT-4 from the source judgements. The higher consistency indicates that the generated meta-reviews follow a more similar sentiment aggregation scheme to get the meta-review sentiments.
Specifically, we first extract judgements from the generated meta-review with a set of Content Expression, $E$ and Sentiment Level, $L$; next, for all expressions in $E$, we get the predicted set of Sentiment Level, $L^{\prime}$ with the zero-shot sentiment prediction function based on judgements in the source documents; last, the accuracy between $L$ and $L^{\prime}$ is the consistency and the final score of FusionEval. FusionEval only considers the precision instead of the recall for meta-review sentiments as it is reference-free and we have no information about the count of judgements that should be synthesized.
@@ 4.2 Measuring Sentiment Fusion for Individual Facets
Table 7: Performances of different LLMs with different prompting methods. (×$\times 0.01$)
@@
Meta-review generated by GPT-4 with the naive prompting
@@
Meta-review generated by GPT-4 with the proposed prompting based on our sentiment consolidation framework
@@ 5 Enhancing LLMs with Explicit Information Consolidation
In this section, to enhance LLMs in meta-review generation we propose two prompting methods to integrate the information consolidation logic. We compare the two methods with other prompting strategies including naive prompting and prompting with LLM-generated logic. We also run experiments on open-source models besides OpenAI ones to investigate the influence of different prompting methods on different models. The experiments are based on automatic and human annotation on 500 samples from PeerSum. [^12] To avoid data contamination, we only use samples which was produced in and after 2022.
@@ 5.1 Prompting LLMs with Sentiment Consolidation Logic
As human meta-reviewers follow the sentiment consolidation logic introduced in Figure 1, LLMs may get improved if we integrate the logic into the generation process. As we know, there are mainly two paradigms in text generation: prompting and fine-tuning. Because fine-tuned models are black-box and it is not easy to integrate the information consolidation logic. Prompting with detailed logic would make it possible to investigate intermediate results in the generation process and this could promote trust and transparency, therefore we focus on enhancing LLMs in prompting with the integration of information logic.
Based on the framework in Figure 1 and comprehensive analysis in Section 3.3, there should be several steps in the meta-review generation process. (1) Step 1: Extracting content and sentiment expressions of judgements in all source documents;(2) Step 2: Predicting Criteria Facets, Sentiment Levels, and Convincingness Levels; (3) Step 3: Reorganize extracted judgements in different clusters for different criteria facets; (4) Step 4: Generate a small summary for judgements on the same criteria facet with sentiment comparison and aggregation; (5) Step 5: Generate the final meta-review based on summaries for different criteria facets.
To integrate the information consolidation logic into prompting of LLMs, we present two methods. (1) Prompt-Ours: we inject descriptive sentiment consolidation logic into the prompt and make the model follow the steps. The complete prompt is in Appendix G.1. (2) Pipeline-Ours: we formulate the task of generating meta-reviews into a pipeline and conduct each part separately instead of a single prompt. The first two steps are finished by GPT-4131313We only use GPT-4 to automate the extraction of judgements, as we find that other models are not well following the instructions. and all used prompts are in Appendix G.2.
@@ 5.2 Reference-Based and Reference-Free Automatic Evaluation
In automatic evaluation, we adopt ROUGE (Lin and Hovy, 2003) [^14] We use the implementation of the algorithm in https://pypi.org/project/rouge-score/ and our proposed FacetEval and FusionEval. Besides ROUGE which is the most popular metric based on surface matching, there are some other semantic representation based metrics like BERTScore (Zhang et al., 2020), UniEval (Zhong et al., 2022), and G-Eval (Liu et al., 2023), but they are only optimized for news summarization or dialogue generation tasks instead of sentiment summarization.
To evaluate the benefit from the integration of the sentiment consolidation logic, we compare the two proposed prompting methods with the naive prompting (Prompt-Naive) which only simply make the model generate meta-reviews without detailed instructions and LLM-generated instructions (Prompt-LLM) which are based on the understanding of the model on how to generate meta-reviews. Moreover, we run experiments on closed-source (GPT-4 and GPT-3.5) and open-source models (LLaMA2-70B and LLaMA2-7B.)151515gpt-4-0613, gpt-3.5-turbo-1106, LLaMA2-70B-Chat, LLaMA2-7B-Chat
Results are present in Table 7. It is clear that all the experimental models perform much better with our proposed two prompting methods, Prompt-Ours and Pipeline-Ours, than other two prompting methods in terms of FusionEval, FacetEval, ROUGE. This show that prompting based on our proposed framework makes the models get similar meta-reviews to the human-written ones and achieve better sentiments in the generated meta-reviews. We find that GPT-4 is still the best model in the scientific sentiment summarization, and the steps generated by the model itself cannot always get better results than the naive prompt. Although Pipeline-Ours can get better results in terms of FusionEval and FacetEval, it performs much worse in terms of ROUGE. This is because it only considers the sentiments rather than other factual summary of the manuscript, which leads to low ROUGE.
@@ 5.3 Case Study on Generated Meta-Reviews
To dive deeper into what difference the integration of sentiment consolidation logic makes, we also conduct a case study on generated results with different prompting methods. We find that generated meta-reviews all seem plausible and machine-generated meta-reviews are much longer than human-written ones and in machine-generated meta-reviews there are many more details which are unnecessary to form an effective meta-review. As shown in the example in Table 8, details about such as "PGD" or "CIFA-10" are not essential to form the meta-review.
We also find that prompts based on our sentiment consolidation framework prefer to cover more balanced judgements based on annotation like in Table 8 and compare them, while the naive prompting only mentions some sentiments. For example, in Table 8, the generation does not include positive sentiments for Clarity and only mentions some issues, but the generation by the promting based on our sentiment consolidation framework acknowledges the good clarity while also mentioning some clarity problems. This is consistent with that prompting with naive prompts get worse sentiments than prompting with the prompt based on the sentiment consolidation logic in Table 7.
@@ 6 Conclusions and Future Work
We initialize the task of scientific sentiment summarization. Based on the nature of the meta-reviewing process, we propose the three-layer framework of sentiment consolidation. With the framework, we propose automatic evaluation metrics to evaluate sentiments in generated summaries. Extensive experiments show that generation results of LLMs could be improved by integrating the sentiment consolidation framework into the textual prompting or the generation pipeline. As the sentiment consolidation also exist in other domains where human reviews or comments exist such as politics and advertisement, we will explore adapting our proposed sentiment consolidation framework into other domains in the future.
@@
Although integration of the sentiment consolidation framework could improve the generation results, there are some limitations of this work:
@@
    We only inject the information consolidation logic into prompting based models instead of fine-tuning based models. We will investigate leveraging the information consolidation logic to improve fine-tuned models in the future.
@@ Appendix B Annotation Instructions for Human Annotation
The screen shots of the two-page annotation instruction for human annotation are shown in Figure 4 and Figure 5.
@@ Appendix C Inter-Annotator Agreement Among Human Annotators and GPT-4
We describe how we calculate inter-annotator agreement among human annotators and GPT-4 here. For Content Expression and Sentiment Expression, as they are highlighted text spans we calculate the character-level agreement with Krippendorf’s $\alpha$ and Cohen’s $\kappa$. Specifically, for each document two annotators may highlight different text spans for Content Expression and Sentiment Expression. We construct two vectors of the same length as the characters to represent highlighting behaviours of any two annotators. This agreement shows whether annotators identify sentiments from similar text spans.
For Criteria Facet, Sentiment Level, and Convincingness Level, we calculate Krippendorf’s $\alpha$ and Cohen’s $\kappa$ in a common way. We first identify whether two annotators recognize sentiment from the same text span with a ROUGE threshold (the summation of ROUGE-1, ROUGE-2 and ROUGE-L between highlighted text spans for a sentiment is larger than 2.0), and calculate agreement on the predicted values.
@@
Annotation Cohen’s $\kappa$ Krippendorf’s $\alpha$ Content Expression 0.623 0.623 Sentiment Expression 0.666 0.665 Criteria Facet 0.769 0.769 Sentiment Level 0.770 0.770 Convincingness Level 0.534 0.533
@@
Annotation Cohen’s $\kappa$ Krippendorff’s $\alpha$ Content Expression 0.631 0.631 Sentiment Expression 0.654 0.654 Criteria Facet 0.783 0.783 Sentiment Level 0.844 0.844 Convincingness Level 0.405 0.398
@@
Annotation Cohen’s $\kappa$ Krippendorff’s $\alpha$ Content Expression 0.572 0.572 Sentiment Expression 0.609 0.609 Criteria Facets 0.857 0.857 Sentiment Levels 0.764 0.763 Convincingness Levels 0.455 0.437
@@
Annotation A B Avg Content Expression 0.558 0.542 0.550 Sentiment Expression 0.565 0.594 0.580 Criteria Facets 0.588 0.610 0.599 Sentiment Levels 0.552 0.541 0.547 Convincingness Levels 0.213 0.192 0.203
@@
Annotation A B Avg Content Expression 0.522 0.534 0.528 Sentiment Expression 0.544 0.569 0.557 Criteria Facets 0.579 0.637 0.608 Sentiment Levels 0.594 0.589 0.592 Convincingness Levels 0.008 0.013 0.011
@@
Annotation A B Avg Content Expression 0.176 0.187 0.182 Sentiment Expression 0.182 0.188 0.185 Criteria Facets 0.480 0.381 0.431 Sentiment Levels 0.123 0.046 0.082 Convincingness Levels 0.0 0.0 0.0

### v2

@@ 4 Sentiment-Aware Evaluation on Information Consolidation
In this section, we focus more on how to evaluate the sentiments of the generated summaries or meta-reviews in meta-review generation based on our proposed framework. We propose FacetEval and FusionEval which are reference-based and reference-free metrics, respectively.
@@ 4.1 Measuring Sentiment Similarity to Human-Written Meta-Review
To assess the quality of generated meta-reviews, we propose a reference-based evaluation metric, FacetEval, measuring the sentiment consistency $c$ between the generated meta-review and the corresponding human-written meta-review in all review facets. Different from the generic evaluation metrics for abstractive summarization or text generation which mostly adopt surface-form matching, we focus more on review facets and their corresponding sentiment levels.
Specifically, we use the distribution of sentiments in all review facets to represent the meta-review and use the cosine similarity of the two vectors as the final score $s$.
@@
where $\bm{d}^{f}=[P^{+}_{f},P_{f},N^{+}_{f},N_{f},O_{f}]$. Results shown in Table 6 suggest that different reviews are consistent in the sentiment to Compliance while there is much lower consistency in Clarity and Novelty. Moreover, we find that conflict reviews [^12] The same as in PeerSum (Li et al., 2023a), if any two reviews have ratings where the gap is larger than 4 they are conflict reviews. would prefer showing conflicts in Advancement, Novelty, Clarity and Overall. This is also consistent with our typical understanding of peer reviews and occasional conflicts among them.
Review Facet w/ Conflicts w/o Conflicts Advancement 0.463 (0.135) 0.551 (0.137) Soundness 0.526 (0.158) 0.501 (0.110) Novelty 0.300 (0.159) 0.357 (0.168) Overall 0.433 (0.147) 0.597 (0.172) Clarity 0.317 (0.133) 0.337 (0.145) Compliance 0.827 (0.071) 0.771 (0.118)
@@
Table 7: Performances of different LLMs with different prompting methods. For all metrics, a larger value denotes better performance. The bold and underlined values are the best and second in each row, respectively (×$\times 0.01$)
Competition Groups Preferred IAA Prompt-Naive LLaMA2-70B 46.67% 0.64 Prompt-Ours LLaMA2-70B 53.33% Prompt-Ours GPT-4 73.33% 0.74 Human-Written 26.67%
Table 8: Two groups of human evaluation results based on human preferences: (1) comparing generated meta-reviews by Prompt-Naive and Prompt-Ours, and (2) comparing human-written meta-reviews and those generated by Prompt-Ours. IAA denotes inter-annotator agreement calculated with nominal Krippendorff’s $\alpha$.
@@
Meta-Review Generated by GPT-4 with Prompt-Naive
@@
Meta-Review Generated by GPT-4 with Prompt-Ours
@@
4.2 Measuring Sentiment Fusion for Individual Facets
Sentiments in the generated meta-reviews should be in line with the aggregate sentiment from the individual source documents including reviews and discussions. Seeing GPT-4 can predict the overall sentiment using judgements from source documents (Section 3.3), we introduce a reference-free evaluation metric, FusionEval, which assesses the consistency between the sentiments in the generated meta-review and that predicted by GPT-4 (with zero-shot prompting) from the source documents. Higher consistency implies the overall sentiment in generated meta-reviews are representative of the sentiments in the reviews and discussions (source documents).
Specifically, we first extract judgements from the generated meta-review following Section 3.2, and these judgements consist of Content Expressions, $E$ and Sentiment Levels, $L$, and the corresponding Review Facets, $F$. Next, for each expression, $e\in E$, we predict the Sentiment Level, $l^{\prime}$, using GPT-4 (zero-shot) based on all judgements for the same Review Facet in the source documents following Section 3.3, and we get predicted Sentiment Levels for all judgements, $L^{\prime}$. Lastly, FusionEval computes an accuracy score by evaluating $L^{\prime}$ against $L$. FusionEval only considers the precision instead of the recall for meta-review sentiments as it is reference-free and we have no information about the count of judgements that should be synthesized.
@@ 5 Enhancing LLMs with Explicit Information Consolidation
In this section, we propose two prompting methods to integrate the sentiment consolidation framework to generate meta-reviews. We compare the two methods with other prompting strategies including naive prompting and prompting with LLM-generated logic. We also run experiments on open-source models besides OpenAI ones to investigate the influence of different prompting methods on different models. The experiments are based on automatic and human annotation on 500 samples from PeerSum. [^13] To avoid data contamination, we only use samples which were produced in and after 2022.
@@ 5.1 Prompting LLMs with Sentiment Consolidation Logic
Following the process in Figure 1 we propose decomposing the meta-review generation process in the following steps: (1) Extracting content and sentiment expressions of judgements from source documents; (2) Predicting Review Facets, Sentiment Levels, and Convincingness Levels; (3) Clustering extracted judgements for different review facets; (4) Generating a “mini summary” for judgements on the same review facet; and (5) Generating the final meta-review based on the mini summaries for all review facets.
We explore two methods to integrate this process for prompting an LLM. (1) Prompt-Ours: we describe the five steps in a single prompt and ask GPT-4 to generate the final meta-reviews (full prompt in Appendix G.1); (2) Pipeline-Ours: we create one prompt for each of the five steps, where the input for the intermediate step is the output from the previous step (full prompts in Appendix G.2).
We experiment with four open-source and close-source LLMs: GPT-4, GPT-3.5, LLaMA2-70B and LLaMA2-7B. [^14] Precise model names for them are: gpt-4-0613, gpt-3.5-turbo-1106, LLaMA2-70B-Chat, LLaMA2-7B-Chat. Note that for Pipeline-Ours, we always use GPT-4 for the first two steps, as we find that the other LLMs perform poorly for these tasks.
5.2 Baselines
As baselines, we include two more methods: (1) Prompt-Naive: which prompts an LLM with a simple instruction to generate the meta-review (full prompt in Appendix G.3); and (2) Prompt-LLM: where we ask an LLM to self-generate the detailed steps for meta-review generation and we include these steps in the final prompt for meta-review generation (full prompt in Appendix G.4).
@@ 5.3 Reference-Based and Reference-Free Automatic Evaluation
For automatic evaluation, we adopt ROUGE (Lin and Hovy, 2003) [^15] We use the implementation of the algorithm in https://pypi.org/project/rouge-score/, FacetEval (Section 4.1) and FusionEval (Section 4.2). [^16] We do not consider other metrics such as BERTScore (Zhang et al., 2020), UniEval (Zhong et al., 2022), or G-Eval (Liu et al., 2023) as they have only been validated for summarization in news or a more general domain.
We present the results in Table 7. Most LLMs perform better with either of our prompting methods (Prompt-Ours and Pipeline-Ours) than the baselines (Prompt-Naive and Prompt-LLM). Comparing between Prompt-Ours and Pipeline-Ours, the former seems to do best for ROGUE while the latter for FusionEval/FacetEval. As FusionEval/FacetEval assesses the accuracy of the overall sentiment specifically, Pipeline-Ours is marginally better when it comes to getting the overall sentiment correct. Comparing different LLMs, GPT-4 seems to work best, although that is mostly true for FusionEval/FacetEval.
5.4 Reference-Free Human Evaluation
To further validate the effectiveness of our prompting methods, we conduct human evaluations to assess the quality of meta-reviews generated by different prompting methods or written by human meta-reviewers. We recruited three volunteer annotators who are senior PhD students familiar with artificial intelligence research and the peer review process. They are asked to select their preferred meta-reviews based on their own understanding of high-quality meta-reviews without knowing the source. [^17] We use majority voting to get the final human preference.
Prompt-Naive vs Prompt-Ours
We randomly select 30 samples and the annotators are asked to compare the generated meta-reviews by Prompt-Naive and Prompt-Ours (using LLaMA2-70B) and select which one is better. Table 8 shows that the meta-reviews generated by Prompt-Ours are selected more by the annotators.
Prompt-Ours vs Human-Written
We repeat the same experiments, but this time comparing meta-reviews generated by Prompt-Ours (GPT-4) vs. written by humans. Looking at Table 8, interestingly Prompt-Ours are much more preferred by the annotators. We suspect this may be because the generated meta-reviews tend to be more consistent in terms of the amount of detail it writes for each review facet, where else there is more variance for the human-written meta-reviews.
@@ 5.5 Case Study on Generated Meta-Reviews
To dive deeper into what difference the integration of sentiment consolidation framework makes, we also conduct a case study on generated meta-reviews with different prompting methods. We find that generated meta-reviews all seem plausible and machine-generated meta-reviews are much longer than human-written ones. In machine-generated meta-reviews, there are more details which are sometimes unnecessary or redundant. As shown in the example in Table 9, details such as "PGD" or "CIFA-10" are not essential to form the meta-review.
Our proposed Prompt-Ours tend to have a more balanced judgements. For example, in Table 9, Prompt-Naive does not talk about the positive aspects for Clarity and only highlights some issues, but Prompt-Ours comments on both the strengths and weaknesses for Clarity. This is consistent with the finding in Table 7 that Prompt-Naive gets worse sentiments than Prompt-Ours.
@@ 6 Conclusions and Future Work
In this paper, we explore sentiment-focused multi-document information consolidation within the task of scientific sentiment summarization. We introduce a three-layer framework of sentiment consolidation to focus on generating meta-reviews and it considers the sentiments for each review facet in the reviews and discussions. We also propose automatic evaluation metrics that assess the overall sentiments in the generated meta-reviews. Experiments on meta-review generation show that prompting LLMs by following the processes in the three-layer framework results in better meta-reviews, providing an empirical validation of our framework for describing the meta-review writing process. As the sentiment consolidation also exist in other domains where human reviews or comments exist such as politics and advertisement, we will explore adapting our proposed sentiment consolidation framework into other domains in the future.
@@
Although integration of the sentiment consolidation framework could improve the generation results, there are still some limitations of this work.
@@
    We only inject the information consolidation logic into prompting based models instead of fine-tuning based models. We will investigate leveraging the information consolidation framework to improve fine-tuned models in the future.
@@
  • $\bullet$
    Meta-review generation is not only about sentiment prediction, future work has to consider more information such as argumentation in source reviews and justification in meta-reviews.
@@
Acknowledgements
We would thank all the anonymous reviewers for their great comments to improve the work. It is impossible to have this work without the volunteer contributions of the annotators on human evaluations.
@@ Appendix B Annotation Instructions for Human Annotation
The screenshots of the two-page annotation instruction for human annotation are shown in Figure 4 and Figure 5 in the last two pages of the Appendix.
@@ Appendix C Inter-Annotator Agreement Among Human Annotators and GPT-4
We describe how we calculate inter-annotator agreement among human annotators and GPT-4 here. For Content Expression and Sentiment Expression, as they are highlighted text spans we calculate the character-level agreement with Krippendorf’s $\alpha$ and Cohen’s $\kappa$. Specifically, for each document, two annotators may highlight different text spans for Content Expression and Sentiment Expression. We construct two vectors of the same length as the characters to represent the highlighting behaviours of any two annotators. This agreement shows whether annotators identify sentiments from similar text spans.
For Review Facet, Sentiment Level, and Convincingness Level, we calculate Krippendorf’s $\alpha$ and Cohen’s $\kappa$ in a common way. We first identify whether two annotators recognize sentiment from the same text span with a ROUGE threshold (the summation of ROUGE-1, ROUGE-2 and ROUGE-L between highlighted text spans for sentiment is larger than 2.0), and calculate agreement on the predicted values.
@@
Annotation Cohen’s $\kappa$ Krippendorf’s $\alpha$ Content Expression 0.623 0.623 Sentiment Expression 0.666 0.665 Review Facet 0.769 0.769 Sentiment Level 0.770 0.770 Convincingness Level 0.534 0.533
@@
Annotation Cohen’s $\kappa$ Krippendorff’s $\alpha$ Content Expression 0.631 0.631 Sentiment Expression 0.654 0.654 Review Facet 0.783 0.783 Sentiment Level 0.844 0.844 Convincingness Level 0.405 0.398
@@
Annotation Cohen’s $\kappa$ Krippendorff’s $\alpha$ Content Expression 0.572 0.572 Sentiment Expression 0.609 0.609 Review Facets 0.857 0.857 Sentiment Levels 0.764 0.763 Convincingness Levels 0.455 0.437
@@
Annotation A B Avg Content Expression 0.558 0.542 0.550 Sentiment Expression 0.565 0.594 0.580 Review Facets 0.588 0.610 0.599 Sentiment Levels 0.552 0.541 0.547 Convincingness Levels 0.213 0.192 0.203
@@
Annotation A B Avg Content Expression 0.522 0.534 0.528 Sentiment Expression 0.544 0.569 0.557 Review Facets 0.579 0.637 0.608 Sentiment Levels 0.594 0.589 0.592 Convincingness Levels 0.008 0.013 0.011
@@
Annotation A B Avg Content Expression 0.176 0.187 0.182 Sentiment Expression 0.182 0.188 0.185 Review Facets 0.480 0.381 0.431 Sentiment Levels 0.123 0.046 0.082 Convincingness Levels 0.0 0.0 0.0